# Feature Re-Embedding: Towards Foundation Model-Level Performance in Computational Pathology

### v3

@@ 4.3.2 Local is more Appropriate than Global
To investigate the role of local self-attention in computational pathology, we replace the naive MSA in the partitioned regions with N-MSA [40]. The results in Table 4 demonstrate that feature re-embedding within local regions, called +N-MSA (local), outperforms the global method under the same N-MSA. This validates the superiority of local self-attention in mining fine-grained features over global ones. The cost of the performance improvement in local self-attention brings a new problem. The local ones suffer from a higher computational burden than the global ones (around 4 more training time). Our proposed RT can alleviate this problem since it employs a more naive MSA, which significantly reduces the computational cost. Another advantage of local self-attention is that it improves the diversity of the re-embedded features, as shown in Figure 5 (c) and (d). Throughout, the local self-attention fashion is far more appropriate for re-embedding in computational pathology, and our proposed RT ensures both good performance and good efficiency in local self-attention.
@@ 4.3.3 Effects of EPEG
We discuss the impact of various positional encoding methods that can handle variable input lengths in detail here. Table 5 shows that the conventional conditional positional encoding methods for the input, such as PEG [5] and PPEG [26], do not effectively improve the performance. More parameters and more complex structures do not bring significant improvements, but rather the simplest PEG achieves slight improvements on both tasks. In contrast, the EPEG, which is embedded in MSA, can benefit from a more lightweight 1-D convolution, and encode the attention matrix more directly. This enables it to model the positional information more effectively in the re-embedding module. For instance, EPEG obtains 0.50%, 0.39%, and 1.74% improvements on C16, NSCLC, and LUAD, respectively.
@@ 4.3.4 Impact of Cross-Region MSA
![[Uncaptioned image]](extracted/5524706/imgs/cw_msa.png)
@@ 4.4 Visualization
We visualize the features before and after re-embedding with different ways in Figure 5. From observations, we can summarize that: 1) The offline feature extractor fails to learn the discriminative instance features, even with a foundation model trained on 200K slide-text pairs, especially when the instance distribution is extremely imbalanced, e.g., 1:224 positive-to-negative ratio; 2) Global self-attention enables the re-embedding of instance features, but its attention distribution is almost uniform. This indicates that its re-embedded features are homogeneous and lack diversity, which limits the performance of advanced MIL models. 3) The features re-embedded by RT not only enhance discriminability but also alleviate the issue of feature homogenization.
@@ D.1 More on Foundation Model Features
In this section, we evaluate the improvement of RT on foundation model features with more MIL models. Figure 6 shows the advantage of RT by online fine-tuning, which can further enhance the discriminability of foundation model features on multiple tasks and models. Surprisingly, we find that this improvement does not decrease with more advanced MIL models, such as RT+CLAM often outperforms RT+ABMIL.
@@ D.3 Region Partition Strategy
| DTFD [41] | ImageNet-1K | 0.79M | 5.1s | 2.1G | 325 | 95.15 | 95.55 | 59.48 |
| TransMIL [22] | ImageNet-1K | 2.67M | 13.2s | 10.6G | 76 | 93.51 | 94.97 | 64.11 |
@@
| ABMIL+N-MSA [36] | ImageNet-1K | 1.64M | 7.7s | 7.2G | 158 | 96.20 | 95.51 | 63.99 |
| RT-MIL(w/o CR-MSA) | ImageNet-1K | 1.64M | 6.1s | 10.0G | 272 | 96.89 | 96.24 | 63.03 |
| RT-MIL | ImageNet-1K | 2.70M | 6.5s | 10.1G | 236 | 97.32 | 96.40 | 67.19 |
| RT-MIL [w/ FFN] | ImageNet-1K | 6.90M | 9.9s | 12.0G | 114 | 96.23 | 95.58 | 64.89 |
| RT-MIL [w/ FFN] | ImageNet-1K | 10.05M | 14.4s | 16.3G | 71 | 96.57 | 95.70 | 63.07 |
@@ D.4 More Parameters are not Always Better
Generally, a Transformer is a multi-layer structure that stacks several blocks with the same structure [21, 33, 17, 9]. However, due to the task specificity, RT only contains a few blocks. Here, we systematically investigate the impact of different numbers of layers and different blocks on the performance and computational cost of RT. First, Figure 9 shows two different blocks: (a) is used by default in RT; (b) introduces a feed-forward network (FFN), which plays an indispensable role in Transformer for NLP or natural image computer vision tasks. From Table 8, we can find that FFN introduces a large number of parameters and computation, but the more expensive computation cost does not bring performance improvement.
Moreover, we can summarize that: 1) Transformer-based methods, represented by TransMIL, bring better long-sequence modeling ability, but also introduce several times more parameters. RT-MIL, as one of them, has equal or less parameter size compared to TransMIL and better performance. 2) As Transformer-based methods, Re-embedding paradigm and RT have Higher parameter efficiency. “+N-MSA” has the same model structure as TransMIL (N-MSA2), but thanks to the Re-embedding, it can achieve higher performance with lower cost. Moreover, RT (w/o CR-MSA) leverages a more excellent design to further improve performance and training time. This indicates that RT-MIL has a good parameter compression space, and can achieve significant improvement with 2 parameters compared to DTFD, reaching the level of the foundation model. 3) In computational pathology, limited by the number of slides, the models face the problem of over-fitting[45], and higher parameter size does not imply better performance. Not only does TransMIL perform poorly on some tasks, we also add FFN and increase the number of layers on the RT-MIL to increase the parameters of RT-MIL (+7.35M), but the performance drops significantly (-4.12% on LUAD).
@@ D.6 Discussion of Hyper-parameter in CR-MSA
The bottom part of Figure 10 shows the results. We can find that RT is not sensitive to this parameter, and different values can achieve high-level performance. In addition, different offline features show similar consistency. This reflects the generality of the preset optimal parameter in different scenarios.

### v4

@@ 4.3.2 Local is more Appropriate than Global
To investigate the role of local self-attention in computational pathology, we replace the naive MSA in the partitioned regions with N-MSA [40]. The results in Table 4 demonstrate that feature re-embedding within local regions, called +N-MSA (local), outperforms the global method under the same N-MSA. This validates the superiority of local self-attention in mining fine-grained features over global ones. The cost of the performance improvement in local self-attention brings a new problem. The local ones suffer from a higher computational burden than the global ones (around 4 more training time). Our proposed R2T can alleviate this problem since it employs a more naive MSA, which significantly reduces the computational cost. Another advantage of local self-attention is that it improves the diversity of the re-embedded features, as shown in Figure 5 (c) and (d). Throughout, the local self-attention fashion is far more appropriate for re-embedding in computational pathology, and our proposed R2T ensures both good performance and good efficiency in local self-attention.
@@ 4.3.3 Effects of EPEG
We discuss the impact of various positional encoding methods that can handle variable input lengths in detail here. Table 5 shows that the conventional conditional positional encoding methods for the input, such as PEG [5] and PPEG [26], do not effectively improve the performance. More parameters and more complex structures do not bring significant improvements, but rather the simplest PEG3×3 achieves slight improvements on both tasks. In contrast, the EPEG, which is embedded in MSA, can benefit from a more lightweight 1-D convolution, and encode the attention matrix more directly. This enables it to model the positional information more effectively in the re-embedding module. For instance, EPEG obtains 0.50%, 0.39%, and 1.74% improvements on C16, NSCLC, and LUAD, respectively.
@@ 4.3.4 Impact of Cross-Region MSA
![[Uncaptioned image]](extracted/5753680/imgs/cw_msa.png)
@@ 4.4 Visualization
We visualize the features before and after re-embedding with different ways in Figure 5. From observations, we can summarize that: 1) The offline feature extractor fails to learn the discriminative instance features, even with a foundation model trained on 200K slide-text pairs, especially when the instance distribution is extremely imbalanced, e.g., 1:224 positive-to-negative ratio; 2) Global self-attention enables the re-embedding of instance features, but its attention distribution is almost uniform. This indicates that its re-embedded features are homogeneous and lack diversity, which limits the performance of advanced MIL models. 3) The features re-embedded by R2T not only enhance discriminability but also alleviate the issue of feature homogenization.
@@ 5 Conclusion
6 Acknowledgement
Reported research is partly supported by the National Natural Science Foundation of China under Grant 62176030 and the Natural Science Foundation of Chongqing under Grant cstc2021jcyj-msxmX0568.
@@ D.1 More on Foundation Model Features
In this section, we evaluate the improvement of R2T on foundation model features with more MIL models. Figure 6 shows the advantage of R2T by online fine-tuning, which can further enhance the discriminability of foundation model features on multiple tasks and models. Surprisingly, we find that this improvement does not decrease with more advanced MIL models, such as R2T+CLAM often outperforms R2T+ABMIL.
@@ D.3 Region Partition Strategy
| DTFD [41] | ImageNet-1K | 0.79M+26M | 5.1s | 2.1G | 325 | 95.15 | 95.55 | 59.48 |
| TransMIL [22] | ImageNet-1K | 2.67M+26M | 13.2s | 10.6G | 76 | 93.51 | 94.97 | 64.11 |
@@
| ABMIL+N-MSA [36] | ImageNet-1K | 1.64M+26M | 7.7s | 7.2G | 158 | 96.20 | 95.51 | 63.99 |
| R2T-MIL(w/o CR-MSA) | ImageNet-1K | 1.64M+26M | 6.1s | 10.0G | 272 | 96.89 | 96.24 | 63.03 |
| R2T-MIL | ImageNet-1K | 2.70M+26M | 6.5s | 10.1G | 236 | 97.32 | 96.40 | 67.19 |
| R2T-MIL [w/ FFN]x2 | ImageNet-1K | 6.90M+26M | 9.9s | 12.0G | 114 | 96.23 | 95.58 | 64.89 |
| R2T-MIL [w/ FFN]x3 | ImageNet-1K | 10.05M+26M | 14.4s | 16.3G | 71 | 96.57 | 95.70 | 63.07 |
@@ D.4 More Parameters are not Always Better
Generally, a Transformer is a multi-layer structure that stacks several blocks with the same structure [21, 33, 17, 9]. However, due to the task specificity, R2T only contains a few blocks. Here, we systematically investigate the impact of different numbers of layers and different blocks on the performance and computational cost of R2T. First, Figure 9 shows two different blocks: (a) is used by default in R2T; (b) introduces a feed-forward network (FFN), which plays an indispensable role in Transformer for NLP or natural image computer vision tasks. From Table 8, we can find that FFN introduces a large number of parameters and computation, but the more expensive computation cost does not bring performance improvement.
Moreover, we can summarize that: 1) Transformer-based methods, represented by TransMIL, bring better long-sequence modeling ability, but also introduce several times more parameters. R2T-MIL, as one of them, has equal or less parameter size compared to TransMIL and better performance. 2) As Transformer-based methods, Re-embedding paradigm and R2T have Higher parameter efficiency. “+N-MSA” has the same model structure as TransMIL (N-MSA2), but thanks to the Re-embedding, it can achieve higher performance with lower cost. Moreover, R2T (w/o CR-MSA) leverages a more excellent design to further improve performance and training time. This indicates that R2T-MIL has a good parameter compression space, and can achieve significant improvement with 2 parameters compared to DTFD, reaching the level of the foundation model. 3) In computational pathology, limited by the number of slides, the models face the problem of over-fitting[45], and higher parameter size does not imply better performance. Not only does TransMIL perform poorly on some tasks, we also add FFN and increase the number of layers on the R2T-MIL to increase the parameters of R2T-MIL (+7.35M), but the performance drops significantly (-4.12% on LUAD).
@@ D.6 Discussion of Hyper-parameter in CR-MSA
The bottom part of Figure 10 shows the results. We can find that R2T is not sensitive to this parameter, and different values can achieve high-level performance. In addition, different offline features show similar consistency. This reflects the generality of the preset optimal parameter in different scenarios.